OpenStudy (anonymous):

The graph shows the progress of two cyclists. Which answer choice shows the rates for each cyclist? A. Cyclist 1: 10 ft/s and Cyclist 2: 12 ft/s B. Cyclist 1: 15 ft/s and Cyclist 2: 12 ft/s C. Cyclist 1: 10 ft/s and Cyclist 2: 15 ft/s D. Cyclist 1: 12 ft/s and Cyclist 2: 10 ft/s

OpenStudy (anonymous):

I think A or C

Mehek (mehek14):

C1: (5, 60) and (10,120) \(\dfrac{120 - 60}{10-5} = \dfrac{60}{5} = 12\)

Mehek (mehek14):

so C1 is 12

OpenStudy (anonymous):

so that means D is the answer

Mehek (mehek14):

C2: (4,40) (10,100) \(\dfrac{100-40}{10-4} =\dfrac{60}{6} = 10\)

Mehek (mehek14):

yup it's D

OpenStudy (anonymous):

thanks

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (anonymous):

I gave you a medal may I have one

OpenStudy (anonymous):

because I guessed correctly

Mehek (mehek14):

lol ok

OpenStudy (anonymous):

thanks can you help me with an other